OpenStudy (ashontae19):

Your club is baking strawberry and apple pies for a bake sale. They need at most 18 pies, and cannot have more than 12 apple pies. Write and graph a system of inequalities to model this system

OpenStudy (ashontae19):

@welshfella

OpenStudy (ashontae19):

can anyone help

OpenStudy (anonymous):

ok

OpenStudy (ashontae19):

@tiny10207

OpenStudy (ashontae19):

thanxs anyone who can

OpenStudy (anonymous):

ok so you divide

OpenStudy (anonymous):

@KaylaRdz0405 @blake57roger plz help her

OpenStudy (anonymous):

1.5

OpenStudy (bossvideogamer21):

so that means 11 apple pies and 7 strawberry pies

OpenStudy (ashontae19):

here my answers

OpenStudy (ashontae19):

@Bossvideogamer21 @LegendaryClasher

OpenStudy (welshfella):

let x be number of strawberry pies and y number of apple pies x + y <= 18 y < 12

OpenStudy (bossvideogamer21):

what

OpenStudy (welshfella):

now draw the graphs of x + y =- 18 and y = 12

OpenStudy (ashontae19):

what about what?

OpenStudy (ashontae19):

so was I right about it being b?

OpenStudy (welshfella):

OK i'll check them out

OpenStudy (ashontae19):

k

OpenStudy (welshfella):

yes my second inequality was not quite correct its y <= 12

OpenStudy (welshfella):

but b is correct

OpenStudy (anonymous):

A medal and i will help

OpenStudy (ashontae19):

thanxs

OpenStudy (ashontae19):

@LegendaryClasher you saying he is wrong?

OpenStudy (anonymous):

yes

OpenStudy (ashontae19):

ok explain

OpenStudy (ashontae19):

@LegendaryClasher

OpenStudy (anonymous):

medal first

OpenStudy (ashontae19):

you need to explain y u are right and i promise you i will give you a medel

OpenStudy (anonymous):

everyone lies

OpenStudy (anonymous):

TO ME

OpenStudy (ashontae19):

@welshfella

OpenStudy (welshfella):

lets have another look I've been wrong before and I'll also sometimes be wrong in the future. I'm no robot.

OpenStudy (welshfella):

the total number of pies must be AT THE MOST 18 that means you have 18 or less pies You cant have any more as an inequlality that is x + y <= 18

OpenStudy (welshfella):

follow?

OpenStudy (ashontae19):

yes

OpenStudy (welshfella):

OK and for apple pies you can only have 12 or less because you cannot have more than 12 th's y <= 12 also x>=0 and y>= 0

OpenStudy (welshfella):